OpenStudy (anonymous):

find dy/dx of r(theta) = 4cos(theta) at theta = pi/3 and pi/2

OpenStudy (anonymous):

r(∅)=4 cos∅ r^' (∅)=4 d/dx cos∅ r^' (∅)=-4 sin∅ r^' (π/3)=-4 sin(π/3)=-4∙[√3)/2]=-2∙√3) r^' (π/2)=-4 sin(π/2)=-4∙1=-4

OpenStudy (anonymous):

isn't that just [dtheta/dx\] and not dy/dx? i am thinking it's (f'theta*sin theta+ fcos theta)/(f'theta*cos theta - f sin theta)

OpenStudy (anonymous):

I was assuming that r(theta)=y and (theta)=x therefore dy/dx=dr/dtheta? Which is basically what I did in my calculations...

OpenStudy (anonymous):

okay is see thanks!

OpenStudy (anonymous):

You're welcome!